Suppose a credit market with a good borrowers and 1 − a bad borrowers. The

good borrowers are all identical, and always repay their loans. Bad borrowers

never repay their loans. Banks issue deposits that pay a real interest rate r1 , and

make loans to borrowers. Banks cannot tell the difference between a good bor-

rower and a bad one. Each borrower has collateral, which is an asset that is worth

A units of future consumption goods in the future period. Determine the interest

rate on loans made by banks.


1
Expert's answer
2021-04-15T10:53:49-0400

"A \lt(1 + r_{1})\times PV"

which means that bad borrowers will not pay back "(1 + r_{1})\times PV," but they will pay A. Since the bank's projected profit from lending is zero in equilibrium, we have "\pi= N \times PV\times (1 + r_{1}) + (1-N)\times A \times PV (1 + r) = 0" . Hence,"r_{1} = 1 + r - (1-N)\times A\times PV \times N-1"

The more collateral, the more repayment the bank will receive if a bad borrower goes bankrupt. Borrowers benefit from having more collateral to secure loans.

If "A\ge PV\times(1 + r_{1})" , then the payment to the bank on a loan to a bad borrower will be "(1 + r_{1})\times PV" , since bad borrowers agree to give "(1 + r_{1})\times PV" to the bank, otherwise the bank will arrest A. Since in equilibrium the bank's profit must be zero, then "\pi=PV\times(1 + r_{1}) - PV\times (1 + r) = 0" , which means "r_{1} = r." If A is significant, then there is normal collateral for leveling problems in the credit market.